A company is required to fence off the area around a robot arm to comply with health and safety law. They have 750m of fencing available. Find the maximum area they can fence off?
1
Expert's answer
2020-05-07T15:50:26-0400

We can see that the maximum perimeter of the area is 750 m.


First, let us consider the rectangular shape with the length "l" and width "w" . Therefore,

"p=2(l+w), \; 750 = 2(l+w), \; l+w = 375." (1)

We should find maximum of area, i.e. maximum of "S(l,w)=l\cdot w." From (1) we write "w=375-l" , so area is "S(l) = l(375-l) = -l^2 + 375l."

Let us maximize this function. We may take the derivative and write

"S'(l) = -2l+375."

"S'(l) = 0" when "l=0.5\cdot375 = 187.5\, m ."

Therefore, "w=375-l = 375-187.5 = 187.5\,m," so the area is a square and the maximum "S" is "187.5^2 = 35156.25\, m^2."


However, according to the Isoperimetric inequality (see https://en.wikipedia.org/wiki/Isoperimetric_inequality#On_a_plane) we can see that circle will be the figure with maximum area.

The perimeter of the circle of radius "R" is "2\pi R", so we can calculate the value of radius:

"R = \dfrac{750}{2\pi} \approx 119.37\,m."

The area will be

"S(R) = \pi R^2 = \pi \cdot 119.37^2 = 44762.33\,m^2."

We may note that circle has significantly larger area.